# G-DexGrasp: Generalizable Dexterous Grasping Synthesis Via Part-Aware Prior Retrieval and Prior-Assisted Generation

### v1

@@ 2.1 Task-Agnostic Dexterous Grasping
Due to the high flexibility of dexterous hands, the primary challenge is to synthesize stable and plausible dexterous grasping solutions [32, 13, 16, 34, 22, 40, 39, 41, 8, 46]. Some works [13, 27, 40, 22] employ optimization algorithms and utilize grasping metrics as objective functions, such as force closure [6] for stability and penetration computation  [6, 40] for collision detection. For example, Obman [13] utilizes GraspIt! simulation [29] to produce feasible and collision-free grasps, while DexGraspNet [40] leverages an efficient differentiable force closure estimator to generate numerous stable grasp results. These approaches are able to produce a large variety of physically feasible grasps, but without considering the reasonable and natural arrangement of dexterous hands, thus not convenient for the subsequent manipulation tasks in real scenarios.
On the other hand, data-driven approaches [11, 7, 24, 5, 41, 19] have become popular, utilizing large annotated datasets to learn to generate plausible and human-like grasps. Many works [35, 16, 41, 20] train Variational Autoencoder networks (VAEs) to learn the grasp distribution and sample diversity grasp solution based on the given condition. Some others [28, 43, 4] utilize the recent diffusion model to sample dexterous grasps via the iterative denoising process. While data-driven approaches are capable of efficiently generating diverse and reasonable grasps, generating high-quality grasps remains a tough task.
@@ 2.2 Task-Oriented Dexterous Grasping
One of the primary advantages of using dexterous hands is the high flexibility to conduct a large variety of manipulation tasks. In addition to the stability and plausibility purpose, more and more research studies task-oriented dexterous grasping, i.e. synthesizing the dexterous grasping that is friendly to the subsequent manipulation tasks.
Some related works [3, 35, 48] classify the task intents into several categories, such as ”use”, ”pass”, etc. They collect large amounts of grasping instances, including the task intent, target object, and dexterous hand pose, and train conditional generative networks to learn the distribution of task-oriented grasps. Moreover, AffordPose [14] collects the grasping instances and the fine-grained annotations, i.e. the part-level affordance labels such as ”twist” and ”handle-grasp”. The affordance performs as a fine-grained classification of the task intent, thus producing more reasonable grasps for different objects and task intents.
The rapid development of natural language processing and the pre-trained large language models enable the understanding and reasoning of task intent in the form of natural language. Task2Grasp [4] utilizes the CLIP text encoder [31] to process the task instruction and trains a diffusion model conditioned on the text encoding and the object point cloud to generate hand poses. By contrast, GraspAsYouSay [42] utilizes a more detailed task instruction that describes the finger arrangement to generate the grasping solution. Instead of training networks from scratch, SemGrasp [20] fine-tunes the pre-trained Multimodal Large Language Model (MLLM), i.e. LLaVA [23], to predict grasp tokens based on the task instruction and object point cloud. The grasp tokens are decoded into dexterous hand parameters with a pre-trained hierarchical VAE network.
@@ 3.1 Overview
Our approach operates in three stages. Specifically, as illustrated in Figure 2, by analyzing the task instruction and target object, we first retrieve the relevant grasping instances and organize the generalizable grasping prior. The grasping prior represents the grasping knowledge for the specific scenario and is composed of the representative hand intrinsics and the contact map distribution, as well as the contact part of the target object. Next, we train a generative part-aware grasping network to estimate a reasonable extrinsic hand parameter conditioned on the specified intrinsics and contact part. Finally, taking the retrieved intrinsics and the generated extrinsics as the initial hand pose, we perform an iterative optimization to refine the dexterous hand with physical constraints and the retrieved grasping prior.
@@ 3.2 Generalizable Grasping Prior Retrieval
Task Instruction Analysis. We first prompt the MLLM to determine the fine-grained grasping arrangement. It takes the task instruction $L$ and the rendered image $I$ of object $O$, and outputs the affordance type $A$ and the name of contact part $P$. The image is rendered from the selected viewpoint to show its shape and appearance as much as possible. In our prompt, we provide a set of pre-defined affordance types as well as their descriptions and restrict the MLLM to select one from them. For example, given the image of a kettle and the instruction ”could you please hold the kettle to pour some tea?”, the MLLM should infer the affordance type ”handle-grasp” and the contact part ”handle”.
@@ 3.4 Prior-Guided Grasping Optimization
where $\hat{\mathcal{V}_{i}}$ and $\mathbf{n}^{\hat{\mathcal{V}}}_{i}$ denote the predicted hand mesh vertex and its corresponding normal vector, $O_{i}$ is the nearest object vertex to $\hat{\mathcal{V}_{i}}$, $c_{revpen}$ is a pre-determined threshold. The third is the human-like grasping constraints, containing $\mathcal{L}_{joint}$ [40] to restrict each angle within the range of human joint and $\mathcal{L}_{spen}$ [51] to avoid self-penetration of the dexterous hand.
@@ 4 Experiments
AffordPose [14] 10.55 1.21 9.93 42.5 1.72
GrabNet [35] 11.07 1.47 9.62 45.5 1.63
GraspTTA [15] 7.98 0.92 7.61 46.5 0.88
Ours 2.94 0.29 4.27 71.6 3.93
Ours 2.04 0.22 4.34 96.30 4.44
Table 1: Quantitative comparison. GrabNet and GraspTTA use BERT to encode task instructions as conditions. Ours uses pre-trained models to infer and localize the contact parts, while Ours uses ground-truth part segmentation with inferred contact part selection. We use red to mark the best results and green the second-best results.
@@ 4.1 Experiment Settings
Dataset. We use the AffordPose dataset [14], which provides the dexterous hand, 3D object shape, and part-level affordance label for each grasp instance. This dataset contains 8 types of hand-object interactions with 13 object categories. Among them, we select five grasping-related interactions, i.e. handle-grip, twist, wrap-around-grip, base-support, trigger-squeeze, to form our training set, which is used for the retrieval and training of the generation network.
We select objects from novel categories to test the generalization on unseen objects. We select a total of 92 objects from 11 categories in the OakInk dataset [48], such as frying pan and bowl; 81 objects from 13 categories in PartNet-Mobility dataset [45], such as suitcase and coffee machine, and 126 objects from 14 categories in 3D-Net dataset [44], such as broom and umbrella, to form our test set.
@@
Implementation. Our experiments invoke GPT-4o as the pre-trained MLLM to infer the affordance type and contact part name of the dexterous grasping. We sample $N=2048$ points from the object surface as the input object points. During training, we use the Adam optimizer with a learning rate of 1e-3 to train the network for 400 epochs. We set the batch size as 90. The loss weights are $\mathcal{\lambda}_{V}=100$, $\mathcal{\lambda}_{R}=5$, $\mathcal{\lambda}_{T}=30$, $\mathcal{\lambda}_{KLD}=10$, $\mathcal{\lambda}_{pen}=1$.

### v2

@@ 2.1 Task-Agnostic Dexterous Grasping
Due to the high flexibility of dexterous hands, the primary challenge is to synthesize stable and plausible dexterous grasping solutions [32, 13, 16, 34, 22, 40, 39, 41, 8, 48]. Some works [13, 27, 40, 22] employ optimization algorithms and utilize grasping metrics as objective functions, such as force closure [6] for stability and penetration computation  [6, 40] for collision detection. For example, Obman [13] utilizes GraspIt!  [29] to produce feasible and collision-free grasps, while DexGraspNet [40] leverages an differentiable force closure estimator to generate stable grasp results. These approaches are able to produce a large variety of physically feasible grasps, but without considering the reasonable and natural arrangement of results, thus not convenient for the subsequent manipulation tasks in real scenarios.
On the other hand, data-driven approaches [11, 7, 24, 5, 41, 19, 47] have become popular, utilizing large annotated datasets to learn to generate plausible and human-like grasps. Many works [35, 16, 41, 20] train Variational Autoencoder networks (VAEs) to learn the grasp distribution and sample diversity grasp solution based on the given condition. Some others [28, 44, 4] utilize the recent diffusion model to sample dexterous grasps via the iterative denoising process. While data-driven approaches are capable of efficiently generating diverse and reasonable grasps, generating high-quality grasps remains a tough task.
@@ 2.2 Task-Oriented Dexterous Grasping
One of the main advantages of dexterous hands is their high flexibility for performing diverse manipulation tasks. Beyond ensuring stability and plausibility, increasing attention has been given to task-oriented dexterous grasping—i.e., generating grasps that facilitate subsequent manipulation.
Some related works [3, 35, 50] classify the task intents into several categories, such as “use", “pass", etc. They collect large amounts of grasping instances, including the task intent, target object, and dexterous hand pose, and train conditional generative networks to learn the distribution of task-oriented grasps. Moreover, AffordPose [14] collects the grasping instances and the fine-grained annotations, i.e. the part-level affordance labels such as “twist" and “handle-grasp". The affordance performs as a fine-grained classification of the task intent, thus producing more reasonable grasps for different objects and task intents.
The rapid development of natural language processing and the pre-trained large language models have enabled understanding and reasoning about task intent from natural language. Task2Grasp [4] uses the CLIP text encoder [31] to process the task instruction and trains a diffusion model conditioned on the text encoding and the object point cloud to generate hand poses. By contrast, GraspAsYouSay [42] utilizes more detailed, finger-level instructions to guide grasp generation. SemGrasp [20] fine-tunes the pre-trained Multimodal Large Language Model (MLLM), i.e. LLaVA [23], to predict grasp tokens based on the task instruction and object point cloud, which are then decoded into hand parameters via a pre-trained hierarchical VAE. The concurrent work AffordDexGrasp [43] generates generalizable affordances to assist grasping.
@@ 3.1 Overview
Our approach consists of three stages. As shown in Figure 2, we first analyze the task instruction and target object to retrieve relevant grasping instances and organize a generalizable grasping prior, which includes representative hand intrinsics, contact map distribution, and the target’s contact part. Next, we train a part-aware generative network to estimate extrinsic hand parameters conditioned on the given intrinsics and contact part. Finally, using the retrieved intrinsics and generated extrinsics as the initial pose, we refine the hand via iterative optimization with physical constraints and the retrieved prior.
@@ 3.2 Generalizable Grasping Prior Retrieval
Task Instruction Analysis. We first prompt the MLLM to determine the fine-grained grasping arrangement. It takes the task instruction $L$ and the rendered image $I$ of object $O$, and outputs the affordance type $A$ and the name of contact part $P$. The image is rendered from the selected viewpoint to show its shape and appearance as much as possible. In our prompt, we provide a set of pre-defined affordance types as well as their descriptions and restrict the MLLM to select one from them. For example, given the image of a kettle and the instruction “could you please hold the kettle to pour some tea?", the MLLM should infer the affordance type “handle-grasp" and the contact part “handle".
@@ 3.4 Prior-Guided Grasping Optimization
where $\hat{\mathcal{V}_{i}}$ and $\mathbf{n}^{\hat{\mathcal{V}}}_{i}$ denote the predicted hand mesh vertex and its normal vector, $O_{i}$ is the nearest object vertex to $\hat{\mathcal{V}_{i}}$, $c_{revpen}$ is a pre-determined threshold. The third is the human-like grasping constraints, containing $\mathcal{L}_{joint}$ [40] to restrict each angle within the range of human joint and $\mathcal{L}_{spen}$ [53] to avoid self-penetration of the hand.
@@ 4 Experiments
AffordPose [14] 6.86/10.55 0.96/1.21 10.66/9.93 81.0/42.5 3.84/1.72
GrabNet [35] 8.76/11.07 1.08/1.47 10.58/9.62 75.6/45.5 3.67/1.63
GraspTTA [15] 5.34/7.98 0.89/0.92 8.17/7.61 74.3/46.5 2.03/0.88
Ours 3.13/2.94 0.82/0.29 5.39/4.27 82.6/71.6 3.72/3.93
Ours 3.18/2.04 0.87/0.22 4.93/4.34 89.74/96.3 3.94/4.44
Table 1: Quantitative comparison on the seen///unseen object categories. GrabNet and GraspTTA use BERT to encode task instructions. Ours uses pre-trained models to infer and localize contact parts. Ours uses ground-truth part segmentation with inferred contact part selection. Best and second-best results are marked in red and green, averaged over seen and unseen cases.
@@ 4.1 Experiment Settings
Dataset. We use the AffordPose dataset [14], which provides dexterous hand data, 3D object shapes, and part-level affordance labels for each grasp. It includes 8 interaction types across 13 object categories. We select five grasp-related interactions—handle-grip, twist, wrap-around-grip, base-support, and trigger-squeeze—and split the data 8:1:1 for training, validation, and testing. The training set is used for retrieval and generation network training, while the test set validates performance on seen categories.
We select novel objects to test generalization, including 92 objects from 11 categories in OakInk [50] (e.g., frying pan, bowl), 81 objects from 13 categories in PartNet-Mobility [46] (e.g., suitcase, coffee machine), and 126 objects from 14 categories in 3D-Net [45] (e.g., broom, umbrella), forming our unseen test set.
@@
Implementation. We invoke GPT-4o as the pre-trained MLLM to infer the affordance type and contact part of the dexterous grasping. We sample $N=2048$ points from the object surface as the input object points. During training, we use the Adam optimizer with a learning rate of 1e-3 to train the network for 400 epochs. The loss weights are $\mathcal{\lambda}_{V}=100$, $\mathcal{\lambda}_{R}=5$, $\mathcal{\lambda}_{T}=30$, $\mathcal{\lambda}_{KLD}=10$, $\mathcal{\lambda}_{pen}=1$.
@@ 5 Conclusion
Acknowledgments.
This work was supported by National Natural Science Foundation of China (62272082, 62302269, 62027826, 62322207), the Excellent Young Scientists Fund Program (Overseas) of Shandong Province (2023HWYQ-034), Shenzhen University Natural Sciences 2035 Program (2022C007), and the joint Fund General Project of Liaoning Provincial Department of Science and Technology (2024-MSLH-352).